Determine the force in each member of the truss shown. State whether each member is in tension or compression and tabulate the results.


1
Expert's answer
2021-04-12T02:31:48-0400


Free body: Entire truss:

+↑ΣFy = 0: By = 0 By = 0

+↺ΣMc = 0: -Bx(3.2 m) - (48 kN)(7.2 m) = 0

Bx = -108 kN Bx = 108 kN ←

+→ΣFx = 0: C - 108 kN + 48 kN = 0

C = 60 kN C = 60 kN →



Free body: Joint B:

FAB/5 = FBC/4 = 108 kN/3

FAB = 180.0 kN T

FBC = 144.0 kN T



Free body: Joint C:

FAC/13 = FBC/12 = 60 kN/5

FAC = 156.0 kN C

FBC = 144 kN (checks)